Company XYZ prepares this for jam, which are labeled to have a mass of 500grams. During the preparation the tins are set to have mean mass of 550grams with standard deviation 20 grams. One tin is selected from a large batch containing 100 and is mass is 500grams. Assuming the masses are normally distributed, test whether the mass of this tin is significantly different from 550 grams at 5% level of significance


1
Expert's answer
2022-01-16T08:52:21-0500

The hypothesis to be tested is given as;

Ho : "\mu" = 550

vs at "\alpha" = 0.05

Ha : "\mu" "\not =" 550

The computed z score is given as follows:


Zo = "\frac{\bar {X}- \mu}{\delta}" , where "\bar {x} = 500, \delta = 20, \mu = 550"

= "\frac{500 -550}{20}"

= "-2.5"

The tabulated z value is;

Z"\alpha/2" = Z0.05/2 = Z0.025 = 1.96 or - 1.96

Since the computed z score = - 2.5 is less than tabulated z value = - 1.96, we reject Ho at 5% level of significance and conclude that the mass of the selected tin is significantly different from 550 grams.